Bond Valuation with Semiannual Payments

Renfro Rentals has issued bonds that have a 6% coupon rate, payable semiannually. The bonds mature in 9 years, have a face value of $1,000, and a yield to maturity of 7%. What is the price of the bonds? Round your answer to the nearest cent.
Answer #1

Price of a bond is mathematically represented as: 1 1 (1+ i) P C i (1) where P is price of bond, with periodic coupon C, n p

M = $1000, n = 9 * 2 = 18 semi-annual periods, i = 7%/2 = 3.5% (semi-annually), C = 6% * $1000/2 = $30 (semi-annually)

P = 30 * \frac{1-\frac{1}{(1 + 0.035)^{18}}}{0.035} + \frac{1000}{(1 + 0.035)^{18}}

P = $395.69 + $538.36

P = $934.1